Factor the expression25a^2-40ab+16b^225a^2-40ab+16b^2 into a product of binomials
Answer by mananth(16949)

25a^2-40ab+16b^2
Ist term and 3rd term are perfect squares
2*sqrt(25a^2)(16b^2)= 40ab
The expression is a perfect square
a^2+2ab+b^2 = (a+b)^2
(5a+4b)^2